OpenStudy (suwhitney):

PLEASE HELP! :( Write (-i)^79/(-3-7i) in standard form.

OpenStudy (suwhitney):

I am supposed to write it in a+bi format... I got an answer but I am not sure if it is right :/

OpenStudy (anonymous):

i=sqrt(-1) i^2 = -1 i^3 = -i i^4 = 1

OpenStudy (anonymous):

\[\frac{ -i }{ -3-7i }\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 58 }+\frac{ 3 }{ 58 }i\]

OpenStudy (anonymous):

IM GONNA CHANGE MY PROF PIC TO SOME SEXY CHICK SO I GET RESPONSES LIKE THESE GAWD DAMN

OpenStudy (anonymous):

hahaha. Sometimes, someone is just being lucky whether what dp you have. :))

OpenStudy (suwhitney):

Jeffery that's the answer my teacher got, for the denominator I got -40 because at the end I got 9-49 :( please explain

OpenStudy (anonymous):

3^2 +7^2

OpenStudy (suwhitney):

Sorry wasnt meant for you Jeffrey..

OpenStudy (suwhitney):

@Vikt0r Haha I do not consider myself "sexy" whatsoever.. but thanks for the compliment.. even though that's not what I'm doing here. I am legitimately trying to get help with my homework douche bag.

OpenStudy (suwhitney):

and I thought you FOILed out the bottom

OpenStudy (anonymous):

@suwhitney didnt say u were sexy dont flatter urself hun i said i was gonna change it to some sexy retricechik but gl with that

OpenStudy (anonymous):

ya, you're going to get, -49 time i^2 so you will get 49

OpenStudy (anonymous):

got it pretty?

OpenStudy (suwhitney):

@Vikt0r And maybe ill change my picture to some guy who thinks he has good looking arms and pose in a dressing room with my cheap retricephone in clothes I can't afford... just a thought. Now be gone imbecile..

OpenStudy (suwhitney):

@Jeffrey_Calderon So how did you get to 58? And thanks handsome ;)

OpenStudy (anonymous):

(-3-7i) times (-3+7i) right? 9-49i^2 = 9+48 = 58

OpenStudy (anonymous):

@suwhitney u got the looking like a male part with that long @ss face and thats my bedroom lol? and thats a otterbox and underneath that otter box is a iphone 4s

OpenStudy (anonymous):

i=sqrt(-1) i^2 = -1 i^3 = -i i^4 = 1

OpenStudy (anonymous):

@suwhitney but gn and umm gl with that schoollll cuz those poles wont make ya money sweetness

OpenStudy (anonymous):

time to repost

OpenStudy (suwhitney):

@Vikt0r Oh, yah because I have little red signs on the back of my closet door too.. And thanks for the luck but I'm already making a perfect GPA. Have fun trolling and making fun of people the rest of your miserable life.. @satellite73 i couldn't agree more.. @Jeffrey_Calderon do you mind if I message you?

OpenStudy (anonymous):

sure:)